OpenStudy (anonymous):

differentiate tan3x using first principle

OpenStudy (dumbcow):

not sure what is meant by first principle? you would use chain rule here where u = 3x du = 3 d/du tan(u) = sec^2 (u) --> d/dx tan(3x) = 3sec^2 (3x)

OpenStudy (nincompoop):

first principle, based on your notes, lecture or textbook

OpenStudy (anonymous):

thanx i kno the answer is 3sec^2(3x). but someone asked to differentiate in frst principle for 6 marks thats y i asked. thsi is supposed to be 1 mark rit?

OpenStudy (anonymous):

i dont kno wat first principle is abt

OpenStudy (alekos):

are you still there?